Please use Java, thank you!

5.

Hashing

1) Insert the keys E X A M Q U S T I O N in that order into an initially empty table of M = 5 lists, using separate chaining. Use the hash function 11 k %

M to transform the kth letter of the alphabet into a table index. Show the hash table after each insertion.

hown in the following table Use A-1, B 2,. as 20 21 22 23 24 25 26 1 2 345 7910 1112 13 14 15 1617 1 19 2 26
Answer #1

EXAMQUSTION

  • insert "E"

index = 11 * k % M = 11 * 5 % 5 = 55 % 5 = 0

index = 0

01234

  • insert "X"

index = 11 * k % M = 11 * 24 % 5 = 264 % 5 = 4

index = 4

  • insert "A"

index = 11 * k % M = 11 * 1 % 5 = 11 % 5 = 1

index = 1

0 소 2-34

  • insert "M"

index = 11 * k % M = 11 * 13 % 5 = 143 % 5 = 3

index = 3

o 2-34

  • insert "Q"

    index = 11 * k % M = 11 * 17 % 5 = 187 % 5 = 2

    index = 2

  • insert "U"

    index = 11 * k % M = 11 * 21 % 5 = 231 % 5 = 1

    index = 1

    0 o 2-34
  • insert "S"

    index = 11 * k % M = 11 * 19 % 5 = 209 % 5 = 4

    index = 4

    ㅇ 소 2-34
  • insert "T"

    index = 11 * k % M = 11 * 20 % 5 = 220 % 5 = 0

    index = 0

    o 2-34
  • insert "I"

    index = 11 * k % M = 11 * 9 % 5 = 99 % 5 = 4

    index = 4

    키| 04 234
  • insert "O"

    index = 11 * k % M = 11 * 15 % 5 = 165 % 5 = 0

    index = 0

    3 M
  • insert "N"

    index = 11 * k % M = 11 * 14 % 5 = 145 % 5 = 4

    index = 4

    3 M